OpenStudy (anonymous):

Teresa is making a small table in the shape of a right triangle. If two of the sides will measure about 14 inches each, how many inches should the longest side (c) measure? Use the appropriate formula and round your answer to the nearest tenth. Answer 22.9 in 19.8 in 16.4 in 18.7 in

OpenStudy (btaylor):

use the pythagorean theorem: you know two sides are 14, so draw the triangle: By the pythagorean theorem, \[c^2= a^2 + b^2 \rightarrow c^2 = 14^2 + 14^2 \rightarrow c^2 = 2(14^2) \rightarrow c = 14\sqrt{2} \rightarrow c \approx 19.8 in\]

OpenStudy (anonymous):

Thanks